# Three quantities arising from Bézout’s identity and resultants of integer polynomials

### v1

@@ 4. Relations between $r(f,g)$ and $R(f,g)$
For any two non-zero coprime integer polynomials $f(x),g(x)$, we have already known that $r(f,g)\mid R(f,g)$ by (1.4). The following result is about the converse of this divisibility relation.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$R(f,g)\mid d^{k}r(f,g)^{\max(n,m)},$
where $n=\deg(f),m=\deg(g),k=\deg(uf)=\deg(vg)$ with polynomials $u,v$ in (3.1).
@@
$\mathrm{Res}(f,r)=r^{\deg(f)}=r^{n}.$
In addition, by (3.1), there exist two integer polynomials $u,v$ such that
$uf+vg=r,$
@@
$\begin{split}\mathrm{Res}(f,r)&=\mathrm{Res}(f,uf+vg)=L(f)^{0-k}\mathrm{Res}(f,vg)\\ &=L(f)^{-k}\mathrm{Res}(f,g)\mathrm{Res}(f,v),\end{split}$
where $k=\deg(vg)=\deg(uf)$. So, we have
$\mathrm{Res}(f,r)=r^{n}=L(f)^{-k}\mathrm{Res}(f,g)\mathrm{Res}(f,v),$
@@
$\mathrm{Res}(g,r)=r^{m}=L(g)^{-k}\mathrm{Res}(g,f)\mathrm{Res}(g,u),$
which implies
(4.2) $L(g)^{k}r^{m}=\mathrm{Res}(g,f)\mathrm{Res}(g,u).$
@@
By the definition of $d$, we write $L(f)=ad,L(g)=bd$ for two coprime positive integers $a,b$ (that is $\gcd(a,b)=1$). Then, combining (4.1) with (4.2), we get
(4.3) $R\mid a^{k}d^{k}r^{n},\qquad R\mid b^{k}d^{k}r^{m}.$
Now, for any prime factor $p$ of $R$, if $p^{j}\mid R$ and $p^{j+1}\nmid R$ for some positive integer $j$, then by (4.3) and noticing $\gcd(a,b)=1$, we must have $p^{j}\mid d^{k}r^{\max(n,m)}$. Hence, we have
$R\mid d^{k}r^{\max(n,m)}.$
This completes the proof. ∎
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$ with $f(x)$ monic, we have $R(f,g)\mid r(f,g)^{n}.$
By Theorem 4.1 and noticing $r(f,g)\mid R(f,g)$ and $d\mid R(f,g)$ (see Theorem 2.1), we directly get that $R(f,g)$ and $dr(f,g)$ have the same prime factors.
@@
For two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d=\gcd(L(f),L(g))$. Then, $R(f,g)$ and $dr(f,g)$ have the same prime factors.
@@ 5. Relations between $B(f,g)$ and $R(f,g)$
Given two non-zero coprime integer polynomials $f(x)$ and $g(x)$, it is known to the experts that $B(f,g)\mid R(f,g)$. We enhance this divisibility relation in the following result.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$dB(f,g)\,\Big{|}\,R(f,g).$
@@
Let $B=B(f,g)$ and $R=R(f,g)$. First, from (1.2) we know that there exist uniquely two integer polynomials $u_{1}(x)$ and $v_{1}(x)$ such that $\deg(u_{1})<\deg(g)$, $\deg(v_{1})<\deg(f)$ and
(5.1) $u_{1}(x)f(x)+v_{1}(x)g(x)=B.$
@@
(5.2) $\gcd(\mathrm{cont}(u_{1}),\mathrm{cont}(v_{1}))=1.$
In addition, by Theorem 2.1 there exist two integer polynomials $u_{2}(x)$ and $v_{2}(x)$ such that $\deg(u_{2})<\deg(g)$, $\deg(v_{2})<\deg(f)$ and:
(5.3) $u_{2}(x)f(x)+v_{2}(x)g(x)=\frac{R}{d}.$
Here, we let $R_{0}=R/d$.
Now, from (5.1) and (5.3), we obtain
$\frac{u_{1}(x)}{B}f(x)+\frac{v_{1}(x)}{B}g(x)=1,$
$\frac{u_{2}(x)}{R_{0}}f(x)+\frac{v_{2}(x)}{R_{0}}g(x)=1.$
By the uniqueness of the polynomials $u(x),v(x)$ in (1.1), we must have
$\frac{u_{1}(x)}{B}=\frac{u_{2}(x)}{R_{0}},\qquad\frac{v_{1}(x)}{B}=\frac{v_{2}(x)}{R_{0}}.$
@@
$B\mathrm{cont}(u_{2})=R_{0}\mathrm{cont}(u_{1}),\qquad B\mathrm{cont}(v_{2})=R_{0}\mathrm{cont}(v_{1}).$
If $B\nmid R_{0}$, then there is a prime number $p$ such that $p\mid B$, $p\mid\mathrm{cont}(u_{1})$ and $p\mid\mathrm{cont}(v_{1})$, and so this contradicts with (5.2). Hence, we must have $B\mid R_{0}$. This completes the proof. ∎
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are not coprime, then $B(f,g)\neq R(f,g)$ and $r(f,g)\neq R(f,g)$.
If the leading coefficients of $f(x)$ and $g(x)$ are coprime, then either $B(f,g)=R(f,g)$ or $B(f,g)\neq R(f,g)$. Both cases can happen; see the following two examples.
@@
Moreover, combining Corollary 3.2 with Corollary 5.2, we directly obtain the following result.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, then $B(f,g)=R(f,g)$ if and only if $r(f,g)=R(f,g)$.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$R(f,g)\mid d^{n+m-1}B(f,g)^{\max(n,m)},$
where $n=\deg(f)$ and $m=\deg(g)$.
@@
Let $B=B(f,g)$ and $R=R(f,g)$. We have known that there exist uniquely two integer polynomials $u_{1}(x)$ and $v_{1}(x)$ such that $\deg(u_{1})<\deg(g)$, $\deg(v_{1})<\deg(f)$ and
$u_{1}(x)f(x)+v_{1}(x)g(x)=B.$
@@
$R\mid d^{k}B^{\max(n,m)},$
where $k=\deg(u_{1}f)=\deg(v_{1}g)\leq n+m-1$. This completes the proof. ∎
By Theorems 5.1 and 5.6, we directly get that $R(f,g)$ and $dB(f,g)$ have the same prime factors.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d=\gcd(L(f),L(g))$. Then, $R(f,g)$ and $dB(f,g)$ have the same prime factors.
@@
Combining Corollary 4.3 with Corollary 5.7, we directly get the following corollary.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are coprime, then $B(f,g)$, $r(f,g)$ and $R(f,g)$ have the same prime factors.
@@ 6. Numerical data
Then, for three positive integers $n,m$ and $H$, we define the following set of pairs of coprime integer polynomials:
$S_{n,m}(H)=$$\Big{\{}(f,g)\in\mathbb{Z}[x]^{2}:\,\deg f=n,\deg g=m,\gcd(f,g)=1,$
$\quad L(f)>0,L(g)>0,\mathrm{H}(f)\leq H,\mathrm{H}(g)\leq H\Big{\}}.$
Note that in the set $S_{n,m}(H)$, both $f$ and $g$ are assumed to have positive leading coefficients, because
@@
In Table 6.1, each value corresponds to $n,m,H$ is the following percentage:
|$\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=r(f,g)\}|}{|S_{n,m}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. Recall that the probability that two randomly given integers are coprime is $6/\pi^{2}\approx 0.6079$; see [18, Theorem 332]. So, by Corollary 3.2, we can roughly say that the above percentage is at least $60\%$ when $H$ is sufficiently large. However, Table 6.1 shows that this percentage is much larger (more than $90\%$). Moreover, when $(n,m)=(1,1)$, this percentages corresponding to $H=30,50$ and $100$ are respectively: $90.29\%$, $90.32\%$, and $90.11\%$.
@@
For any positive integers $n$ and $m$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=r(f,g)\}|}{|S_{n,m}(H)|}\geq\frac{9}{10}.$
$(n,m)$ $H$ 2 3 4 5 6
@@
In Table 6.2, each value corresponds to $n,m,H$ is the following percentage:
|$\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=R(f,g)\}|}{|S_{n,m}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. As the above, by Corollary 5.2 we can roughly say that this percentage is at most $61\%$ when $H$ is sufficiently large. However, Table 6.2 shows that this percentage is much smaller except the case when $(n,m)=(1,1)$. Table 6.2 suggests that for two coprime integer polynomials $f(x),g(x)$ given randomly, we often have $B(f,g)\neq R(f,g)$ when $\deg f$ and $\deg g$ are large.
Regarding Table 6.2, we pose the following conjecture.
@@
For any positive integers $n$ and $m$ with $(n,m)\neq(1,1)$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)\neq R(f,g)\}|}{|S_{n,m}(H)|}>\frac{1}{2}.$
Given two coprime integer polynomials of degree one $f(x)=ax+b$ and $g(x)=cx+d$, we directly have $R(f,g)=|ad-bc|$ and
@@
So, by definition we have $B(f,g)=|ad-bc|/\gcd(a,c)$. Hence, $B(f,g)=R(f,g)$ whenever $\gcd(a,c)=1$. This can explain why the percentages in Table 6.2 corresponding to the case when $(n,m)=(1,1)$ are large.
$(n,m)$ $H$ 2 3 4 5 6

### v2

@@ 4. Relations between $r(f,g)$ and $R(f,g)$
We have already known that $r(f,g)\mid R(f,g)$ by (1.3). The following result is about the converse of this divisibility relation. Recall that $d=\gcd(L(f),L(g))$.
@@
We have
$R(f,g)\mid d^{j}r(f,g)^{\max(m,n)},$
where $m=\deg(f),n=\deg(g),j=\deg(pf)=\deg(qg)$ with polynomials $p,q$ in (3.1). Moreover, if both $f$ and $g$ are monic, then we have
$R(f,g)\mid r(f,g)^{\min(m,n)}.$
@@
$\mathrm{Res}(f,r)=r^{\deg(f)}=r^{m}.$
In addition, by (3.1), there exist two integer polynomials $p,q$ such that
$pf+qg=r,$
@@
$\begin{split}\mathrm{Res}(f,r)&=\mathrm{Res}(f,pf+qg)=L(f)^{0-j}\mathrm{Res}(f,qg)\\ &=L(f)^{-j}\mathrm{Res}(f,g)\mathrm{Res}(f,q),\end{split}$
where $j=\deg(qg)=\deg(pf)$. So, we have
$\mathrm{Res}(f,r)=r^{m}=L(f)^{-j}\mathrm{Res}(f,g)\mathrm{Res}(f,q),$
@@
(4.2) $L(g)^{j}r^{n}=\mathrm{Res}(g,f)\mathrm{Res}(g,p).$
@@
In view of $d=\gcd(L(f),L(g))$, we write $L(f)=ad,L(g)=bd$ for two coprime integers $a,b$ (that is $\gcd(a,b)=1$). Then, combining (4.1) with (4.2), we get
(4.3) $R\mid a^{j}d^{j}r^{m},\qquad R\mid b^{j}d^{j}r^{n}.$
Now, for any prime factor $\ell$ of $R$, if $\ell^{i}\mid R$ and $\ell^{i+1}\nmid R$ for some positive integer $i$, then by (4.3) and noticing $\gcd(a,b)=1$, we must have $\ell^{i}\mid d^{j}r^{\max(m,n)}$. Hence, we have
$R\mid d^{j}r^{\max(m,n)}.$
This gives the first part of the theorem.
Moreover, if both $f$ and $g$ are monic, then by (4.3) we have
$R\mid r^{m}\quad\textrm{and}\quad R\mid r^{n},$
and so, $R\mid r^{\min(m,n)}$. This completes the proof. ∎
We remark that for the general case in Theorem 4.1, the power $\max(m,n)$ can not be replaced by $\min(m,n)$; see Example 4.2 (for $d=1$) and Example 4.3 (for $d>1$).
Example 4.2.
Let $f(x)=2x^{3}+3x^{2}-2$ and g(x)=3x3g(x)=3x-3. Then, $d=1$, $B(f,g)=r(f,g)=9$, and $R(f,g)=81$.
Example 4.3.
Let $f(x)=2x^{3}+x^{2}-3x+2$ and g(x)=4x2g(x)=4x-2. Then, $d=2$, $B(f,g)=r(f,g)=2$, and $R(f,g)=64$. Moreover, in this case we can choose $j=3$, since
$2f(x)-(x^{2}+x-1)g(x)=2.$
@@
If moreover $f(x)$ is monic, then we have $R(f,g)\mid r(f,g)^{m}$, where $m=\deg(f)$.
By Theorem 4.1 and noticing $r(f,g)\mid R(f,g)$ and $d\mid R(f,g)$ (see Theorem 2.1), we directly get that $R(f,g)$ and $dr(f,g)$ have the same prime factors. This result is known to the experts.
@@
$R(f,g)$ and $dr(f,g)$ have the same prime factors.
@@ 5. Relations between $B(f,g)$ and $R(f,g)$
It is known to the experts that $B(f,g)\mid R(f,g)$. We enhance this divisibility relation in the following result. Recall that $d=\gcd(L(f),L(g))$.
@@
We have
$dB(f,g)\mid R(f,g).$
@@
Let $B=B(f,g)$ and $R=R(f,g)$. First, from (1.2) we know that there exist uniquely two integer polynomials $p_{1}(x)$ and $q_{1}(x)$ such that $\deg(p_{1})<\deg(g)$, $\deg(q_{1})<\deg(f)$ and
(5.1) $p_{1}(x)f(x)+q_{1}(x)g(x)=B.$
@@
(5.2) $\gcd(\mathrm{cont}(p_{1}),\mathrm{cont}(q_{1}))=1.$
In addition, by Theorem 2.1 there exist uniquely two integer polynomials $p_{2}(x)$ and $q_{2}(x)$ such that $\deg(p_{2})<\deg(g)$, $\deg(q_{2})<\deg(f)$ and
(5.3) $p_{2}(x)f(x)+q_{2}(x)g(x)=\frac{R}{d}.$
Let $R_{0}=R/d$. Then, from (5.1) and (5.3), we obtain
$\frac{p_{1}(x)}{B}f(x)+\frac{q_{1}(x)}{B}g(x)=1,$
$\frac{p_{2}(x)}{R_{0}}f(x)+\frac{q_{2}(x)}{R_{0}}g(x)=1.$
By the uniqueness of the polynomials $p(x),q(x)$ in (1.1), we must have
$\frac{p_{1}(x)}{B}=\frac{p_{2}(x)}{R_{0}}\quad\textrm{and}\quad\frac{q_{1}(x)}{B}=\frac{q_{2}(x)}{R_{0}}.$
@@
$B\mathrm{cont}(p_{2})=R_{0}\mathrm{cont}(p_{1}),\qquad B\mathrm{cont}(q_{2})=R_{0}\mathrm{cont}(q_{1}).$
If $B\nmid R_{0}$, then there is a prime number $\ell$ such that $\ell\mid B$, $\ell\mid\mathrm{cont}(p_{1})$ and $\ell\mid\mathrm{cont}(q_{1})$, and so this contradicts with (5.2). Hence, we must have $B\mid R_{0}$. This completes the proof. ∎
@@
If $d>1$, then $B(f,g)\neq R(f,g)$ and $r(f,g)\neq R(f,g)$.
If $d=1$, then either $B(f,g)=R(f,g)$ or $B(f,g)\neq R(f,g)$. Both cases can happen; see the following two examples.
@@
Moreover, combining Corollary 3.2 with Corollary 5.2 and noticing $r(f,g)\leq B(f,g)\leq R(f,g)$, we directly obtain the following result.
@@
$B(f,g)=R(f,g)$ if and only if $r(f,g)=R(f,g)$.
@@
We have
$R(f,g)\mid d^{m+n-1}B(f,g)^{\max(m,n)},$
where $m=\deg(f)$ and $n=\deg(g)$. Moreover, if both $f$ and $g$ are monic, then we have
$R(f,g)\mid B(f,g)^{\min(m,n)}.$
@@
Let $B=B(f,g)$ and $R=R(f,g)$. We have known that there exist uniquely two integer polynomials $p_{1}(x)$ and $q_{1}(x)$ such that $\deg(p_{1})<\deg(g)$, $\deg(q_{1})<\deg(f)$ and
$p_{1}(x)f(x)+q_{1}(x)g(x)=B.$
@@
$R\mid d^{j}B^{\max(m,n)},$
where $j=\deg(p_{1}f)=\deg(q_{1}g)\leq m+n-1$. Also, when both $f$ and $g$ are monic, we have
$R\mid B(f,g)^{\min(m,n)}.$
This completes the proof. ∎
We remark that for the general case in Theorem 5.6, the power $\max(m,n)$ can not be replaced by $\min(m,n)$; see Example 4.2 (for $d=1$) and Example 4.3 (for $d>1$).
By Theorems 5.1 and 5.6, we directly get that $R(f,g)$ and $dB(f,g)$ have the same prime factors. This also can be seen from Corollary 4.5.
@@
$R(f,g)$ and $dB(f,g)$ have the same prime factors.
@@
Combining Corollary 4.5 with Corollary 5.7, we directly get the following corollary. We also believe that this result is known to the experts, but we haven’t found a specific reference for it.
@@
If $d=1$, then $B(f,g)$, $r(f,g)$ and $R(f,g)$ have the same prime factors.
@@ 6. Numerical data
Then, for three positive integers $m,n$ and $H$, we define the following set of pairs of coprime integer polynomials:
$S_{m,n}(H)=$ $\Big\{(f,g)\in\mathbb{Z}[x]^{2}:\,\deg(f)=m,\deg(g)=n,L(f)>0,$
$\quad L(g)>0,\textrm{$f$ and $g$ are coprime},\mathrm{H}(f)\leq H,\mathrm{H}(g)\leq H\Big\}.$
Note that in the set $S_{m,n}(H)$, both $f$ and $g$ are assumed to have positive leading coefficients, because
@@
In Table 6.1, each value corresponds to $m,n,H$ is the following percentage:
$\frac{|\{(f,g)\in S_{m,n}(H):\,B(f,g)=r(f,g)\}|}{|S_{m,n}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. Recall that the probability that two randomly given integers are coprime is $6/\pi^{2}\approx 0.6079$; see [23, Theorem 332]. So, by Corollary 3.2, we can roughly say that the above percentage is at least $60\%$ when $H$ is sufficiently large. However, Table 6.1 shows that this percentage is much larger (more than $90\%$). Moreover, when $(m,n)=(1,1)$, this percentages corresponding to $H=30,50$ and $100$ are respectively: $90.29\%$, $90.32\%$, and $90.11\%$.
@@
For any positive integers $m$ and $n$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{m,n}(H):\,B(f,g)=r(f,g)\}|}{|S_{m,n}(H)|}\geq\frac{9}{10}.$
$(m,n)$ $H$ 2 3 4 5 6
@@
In Table 6.2, each value corresponds to $m,n,H$ is the following percentage:
$\frac{|\{(f,g)\in S_{m,n}(H):\,B(f,g)=R(f,g)\}|}{|S_{m,n}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. As the above, by Corollary 5.2 we can roughly say that this percentage is at most $61\%$ when $H$ is sufficiently large. However, Table 6.2 shows that this percentage is much smaller except the case when $(m,n)=(1,1)$.
Table 6.2 suggests that for two coprime integer polynomials $f(x),g(x)$ given randomly, we often have $B(f,g)\neq R(f,g)$ when $\deg(f)$ and $\deg(g)$ are large. Regarding Table 6.2, we pose the following conjecture.
@@
For any positive integers $m$ and $n$ with $m\geq 2$ and $n\geq 2$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{m,n}(H):\,B(f,g)=R(f,g)\}|}{|S_{m,n}(H)|}<\frac{1}{2}.$
Given two coprime integer polynomials of degree one $f(x)=ax+b$ and $g(x)=cx+d$, we directly have R(f,g)=|adbc|R(f,g)=|ad-bc| and
@@
So, by definition we have $B(f,g)=|ad-bc|/\gcd(a,c)$. Hence, $B(f,g)=R(f,g)$ whenever $\gcd(a,c)=1$. This can explain why the percentages in Table 6.2 corresponding to the case when $(m,n)=(1,1)$ are large.
$(m,n)$ $H$ 2 3 4 5 6
@@
7. Comments
In this section, we want to say some more words about the resultant and relevant concepts.
The resultant of two univariate polynomials over the integers $\mathbb{Z}$ is an important concept and useful in algebra and number theory. It has been generalized to a more general setting (see [3, Section 4.2] or [26, Chapter 7]) and also to multivariate polynomials (see [9, Chapter 3, Section 6]). One can also interpret it from the viewpoint of algebraic geometry; see [20, Section V.2]. From this viewpoint, one may define new integers with respect to two coprime integer polynomials.
Moreover, the subresultants of two uivariate integer polynomials are defined by means of submatrices of their Sylvester matrix (see [3, Section 4.2] or [26, Chapter 7] for a more general setting). The subresultants can be viewed as a generalization of the resultant, and they are very useful for computing the resultant and the greatest common divisor of those two polynomials.
In addition, since the polynomial ring $\mathbb{Z}[x]$ is a unique factorization domain, for the two coprime polynomials $f,g\in\mathbb{Z}[x]$ given above, their greatest common divisor $\gcd(f,g)$ exists in $\mathbb{Z}[x]$ and in fact equals to $\gcd(\mathrm{cont}(f),\mathrm{cont}(g))$ (that is, the greatest common divisor of all the coefficients of $f$ and $g$). So, the integer $\gcd(f,g)$ is usually different from the three integers $B(f,g)$, $r(f,g)$ and $R(f,g)$ we have studied.